doglover:

A line contains the points (–5,0) and (0,–3). Challenge: What is the equation, in slope-intercept form, of the line that contains both of these points? 𝑦=_____ and x-____ Explanation: Explain how you found the equation above. "I used the points to find the slope, which is_____" can someone please help me

Mercury:

slope formula m = \[\frac{ x_{2}-x_{1} }{ y_{2}-y_{1} }\] pick (–5,0) to be your (x1,y1) and (0,-3) to be your (x2,y2). with this setup, x1 = -5, y1 = 0, and x2 = 0, y2 = -3 once you have your slope, you can set up the equation y = mx + b plug in your m-value from before, plug in one of your (x,y) coordinates (would recommend (0,-3)), and solve for b once you have both m and b, simply set up y = mx + b and plug your m and b values in. the result is your equation.

doglover:

ok, thank you!